OpenStudy (anonymous):

Hi, I am working on some serious physics and need help to understand c2 = (fy)2 = F / md

OpenStudy (anonymous):

I don't recognize the equation(s). Could you identify the symbols.

OpenStudy (anonymous):

http://zhidao.baidu.com/question/412047513.html?seed=0 comes form this article

OpenStudy (anonymous):

I would really think being able to deduce F=ma E=mc^2 and E=hf in one formulae would give me a decent basic understanding of physics

OpenStudy (anonymous):

I just couldn't find any other webpages with info on this, so I kind of grew suspicious of it being fraud.

OpenStudy (anonymous):

It's related to Milo Wolff

OpenStudy (anonymous):

http://www.quantummatter.com/space-resonance/

OpenStudy (anonymous):

My first observation is that the equation c^2=F/m is incorrect invalidating the exercise. Why?

OpenStudy (anonymous):

why does it invalidate it? are also E = kwy2 also wrong?

OpenStudy (anonymous):

btw do you know anytinhg about this formulae for waves? c2 = (fy)2 = F / mdc = velocity of light f = frequency y = wavelength F = Force = constant (ka) by amplitude (x) = kax md = mass density = mass (m) / wavelength (y) = m / y Let us assume that amplitude is proportional to wavelength F = kax = kwy Energy (E) = force (kwy) by distance (y) thus; E = kwy2

OpenStudy (anonymous):

oh, you mean c^2=F/m gives false results

OpenStudy (anonymous):

c^2=F/M is invalid therefore cannot be used . It will give you false results. note that mc^2= Energy not Force.